### v3

# Generalized positive scalar curvature on spinc manifolds

Abstract.
Let $(M,L)$ be a (compact) non-spin spinc manifold. Fix a Riemannian metric $g$ on $M$ and a connection $A$ on $L$, and let $D_{L}$ be the associated spinc Dirac operator. Let $R^{\text{{tw}}}_{(g,A)}:=R_{g}+2ic(\Omega)$ be the twisted scalar curvature (which takes values in the endomorphisms of the spinor bundle), where $R_{g}$ is the scalar curvature of $g$ and $2ic(\Omega)$ comes from the curvature $2$-form $\Omega$ of the connection $A$. Then the Lichnerowicz-Schrödinger formula for the square of the Dirac operator takes the form $D_{L}^{2}=\nabla^{*}\nabla+\frac{1}{4}R^{\text{{tw}}}_{(g,A)}$. In a previous work we proved that a closed non-spin simply-connected spinc-manifold $(M,L)$ of dimension $n\geq 5$ admits a pair $(g,A)$ such that $R^{\text{{tw}}}_{(g,A)}>0$ if and only if the index $\alpha^{c}(M,L):=\operatorname{ind}D_{L}$ vanishes in $K_{n}$. In this paper we introduce a scalar-valued generalized scalar curvature $R^{\text{{gen}}}_{(g,A)}:=R_{g}-2|\Omega|_{op}$, where $|\Omega|_{op}$ is the pointwise operator norm of Clifford multiplication $c(\Omega)$, acting on spinors. We show that the positivity condition on the operator $R^{\text{{tw}}}_{(g,A)}$ is equivalent to the positivity of the scalar function $R^{\text{{gen}}}_{(g,A)}$. We prove a corresponding trichotomy theorem concerning the curvature $R^{\text{{gen}}}_{(g,A)}$, and study its implications. We also show that the space $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ of pairs $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}>0$ has non-trivial topology, and address a conjecture about non-triviality of the “index difference” map.
Key words and phrases:
positive scalar curvature, spinc manifold, bordism, $K$-theory, index, index difference
2020 Mathematics Subject Classification:
Primary 53C21; Secondary 53C27, 58J22, 55N22, 19L41
BB was partially supported by Simons collaboration grant 708183
1. Introduction
1.1. Motivation and results
Recall that a spinc manifold $(M,L)$ is a smooth oriented manifold $M$, $\dim M=n$, equipped with a choice of a complex line bundle $L$ with the property that $c_{1}(L)$ reduces mod $2$ to $w_{2}(M)$. Then $M$ admits a spinc Dirac operator $D_{L}$, which can be viewed as the spin Dirac operator $D$ on $M$ (which always exists locally, but won’t exist globally on $M$ if $w_{2}(M)\neq 0$) twisted by the “line bundle” $L^{1/2}$. (Note that $L^{1/2}$ is also not globally well-defined, but the indeterminacies in $D$ and in $L^{1/2}$ cancel each other out.) To define the Dirac operator $D_{L}$, one needs some extra data: a Riemannian metric $g$ on $M$, a hermitian bundle metric $h$ on $L$, and a connection $A$ on $L$. The $K$-homology class $[D_{L}]\in K_{n}(M)$ of the operator does not depend on any of these choices, nor does the index $\alpha^{c}(M,L):=\operatorname{ind}D_{L}=p_{*}([D_{L}])\in K_{n}$, where $p\colon\,M\to\text{{pt}}$. The Lichnerowicz-Schrödinger formula for the square of the Dirac operator takes the form
(1.1) $D_{L}^{2}=\nabla^{*}\nabla+\frac{R_{g}}{4}+\mathcal{R}_{L},$
where $R_{g}$ is the scalar curvature of $M$ and $\mathcal{R}_{L}=\frac{1}{2}ic(\Omega)$ is $\frac{i}{2}$ times Clifford multiplication by the curvature $2$-form $\Omega$ of the connection $A$ acting on spinors; see [19, Theorem D.12].
We define the twisted scalar curvature $R^{\text{{tw}}}_{(g,A)}$ to be the function $R^{\text{{tw}}}_{(g,A)}:=R_{g}+2ic(\Omega)$, taking values in sections of the endomorphism bundle of the spinor bundle, so that $D_{L}^{2}=\nabla^{*}\nabla+\frac{1}{4}R^{\text{{tw}}}_{(g,A)}$. This function $R^{\text{{tw}}}_{(g,A)}$ (except for a factor of $4$) appeared in [3], where it was called the curvature endomorphism $\mathcal{K}$ for the spinc spinor bundle. So when $n$ is even, the condition that the operator $R^{\text{{tw}}}_{(g,A)}$ is positive-definite forces $\alpha^{c}(M,L)=\operatorname{ind}D_{L}$ to vanish. We use the notation $R^{\text{{tw}}}_{(g,A)}>0$ and call this condition twisted positive scalar curvature or twisted psc for short.
Note that the condition $\alpha^{c}(M,L)=0$ depends very much on the Chern class $c_{1}(L)$. For example, when $M=\mathbb{C}\mathbb{P}^{2}$, which is spinc but not spin, $c_{1}(L)\in H^{2}(\mathbb{C}\mathbb{P}^{2};\mathbb{Z})\cong\mathbb{Z}$ must be odd since it must reduce mod $2$ to $w_{2}(M)\neq 0$, and the condition $\alpha^{c}(M,L)=0$ is satisfied when $c_{1}(L)=\pm 1$ but not when $|c_{1}(L)|\geq 3$. (See for example [12, Theorem 3.13].)
In our previous paper [5], we proved a surgery theorem for the twisted psc condition:
Theorem 1.1 ([5, Theorem 4.2]).
Let $(M,L)$ be a closed spinc manifold. Assume that $M$ admits a Riemannian metric $g$ and the line bundle $L$ admits a hermitian bundle metric $h$ and a unitary connection $A$ such that $R^{\text{{tw}}}_{(g,A)}>0$. Let $(M^{\prime},L^{\prime})$ be obtained from $(M,L)$ by spinc surgery in codimension $\geq 3$. Then $(M^{\prime},L^{\prime})$ admits a hermitian bundle metric $h^{\prime}$, a Riemannian metric $g^{\prime}$, and a unitary connection $A^{\prime}$ such that $R^{\text{{tw}}}_{(g^{\prime},A^{\prime})}>0$.
Then from this and a bordism argument, we proved the following:
Theorem 1.2 (see [5, Theorem 4.2] and [5, Corollary 5.2]).
Let $(M,L)$ be a simply connected closed spinc manifold which is not spin, with $n=\dim M\geq 5$. Then $(M,L)$ admits a Riemannian metric $g$, a hermitian bundle metric $h$, and a unitary connection $A$ such that $R^{\text{{tw}}}_{(g,A)}>0$, if and only if the index $\alpha^{c}(M,L)$ vanishes in $K_{n}$.
Given a spinc manifold $(M,L)$ we define the generalized scalar curvature $R^{\text{{gen}}}_{(g,A)}$ of a triple $(g,h,A)$, where $g$ is a Riemannian metric on $M$, $A$ is a connection on the line bundle $L$, and $h$ is a metric on the bundle $L$, to be the scalar-valued function $R^{\text{{gen}}}_{(g,A)}:=R_{g}-2|\Omega|_{op}$, where $|\Omega|_{op}$ is the pointwise operator norm of Clifford multiplication $c(\Omega)$, acting on spinors. [^1] Recall that if $x\in M$, $|\Omega|_{op}(x)=\max\{|c(\Omega)\psi(x)|:|\psi(x)|=1\}$, where the maximum is taken over all spinors $\psi$ with norm $1$ at the point $x$. Note that when $M$ is spin, one can choose $L$ to be the trivial bundle and the connection $A$ to be flat, so in this case $R^{\text{{gen}}}_{(g,A)}=R_{g}$. So in that sense $R^{\text{{gen}}}_{(g,A)}$ generalizes the notion of scalar curvature when we pass from spin manifolds to spinc manifolds.
We observe that the positivity condition on the operator $R^{\text{{tw}}}_{(g,A)}$ is equivalent to the positivity of the scalar function $R^{\text{{gen}}}_{(g,A)}$ (see Lemma 3.1).
Recall the trichotomy theorem of Kazdan-Warner, which can be formulated as follows:
Theorem 1.3 (Kazdan-Warner [18, 17, 15]).
Every closed connected manifold $M$ of dimension $\geq 3$ falls into exactly one of the following three classes:
  1. (1)
    those admitting a psc metric, in which case every smooth function on $M$ is the scalar curvature of some metric;
  2. (2)
    those admitting a metric $g$ with $R_{g}\equiv 0$ but not a metric with $R_{g}\geq 0$ and $R_{g}$ positive somewhere — in this case a metric with $R_{g}\equiv 0$ is necessarily Ricci-flat, and the possible scalar curvature functions on $M$ are 00 and the functions negative somewhere;
  3. (3)
    all other manifolds, those not admitting any metric with nonnegative scalar curvature — in this case, the possible scalar curvature functions of metrics on $M$ are exactly those functions which are negative somewhere.
We prove a corresponding trichotomy theorem concerning the generalized scalar curvature $R^{\text{{gen}}}_{(g,A)}$:
Theorem 1.4.
Every closed connected spinc manifold $(M,L)$ of dimension at least three falls into exactly one of the following three classes:
  1. (1)
    those admitting a pair $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}>0$, in which case every smooth function on $M$ is the generalized scalar curvature of some pair $(g^{\prime},A^{\prime})$, $g^{\prime}$ a Riemannian metric on $M$ and $A^{\prime}$ a connection on $L$;
  2. (2)
    those admitting a pair $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}\equiv 0$ but not a pair $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}$ positive everywhere — in this case a smooth function on $M$ is the generalized scalar curvature of some pair $(g^{\prime},A^{\prime})$ if and only if it is either $\equiv 0$ or negative somewhere on $M$;
  3. (3)
    all other manifolds, those not admitting any $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}\geq 0$ — in this case, a smooth function on $M$ is the generalized scalar curvature of some pair $(g^{\prime},A^{\prime})$ if and only if it is negative somewhere on $M$.
Recall that the conformal Laplacian $\mathsf{L}_{g}=-\frac{4(n-1)}{n-2}\Delta_{g}+R_{g}$ and its lowest eigenvalue $\mu_{1}(\mathsf{L}_{g})$ played a major role in proving Theorem 1.3, where the sign of $\mu_{1}(\mathsf{L}_{g})$ (which is invariant under a conformal change of the metric $g$) points to the corresponding class (1), (2) or (3). [^2] We use the same sign convention as Kazdan and Warner, that $\Delta_{g}$ is a nonpositive operator. This explains the initial minus sign in the formula for $\mathsf{L}_{g}$. Similarly, the generalized scalar curvature $R^{\text{{gen}}}_{(g,A)}$ is closely related to the principal eigenvalue $\mu_{1}(\mathsf{L}_{(g,A)})$ of the generalized conformal Laplacian
${\mathsf{L}_{(g,A)}}=-\frac{4(n-1)}{n-2}\Delta_{g}+{R^{\text{{gen}}}_{(g,A)}}=-\frac{4(n-1)}{n-2}\Delta_{g}+R_{g}-2|\Omega|_{op}=\mathsf{L}_{g}-2|\Omega|_{op}.$
Here again, the sign of $\mu_{1}(\mathsf{L}_{(g,A)})$ is invariant under a conformal change of the metric $g$. The condition $\mu_{1}(\mathsf{L}_{(g,A)})>0$ corresponds to the case (1): then there exists a pair $(g^{\prime},A)$ (where $g^{\prime}$ is conformal to $g$) such that $R^{\text{{gen}}}_{(g^{\prime},A)}>0$. In turn, this means that $\alpha^{c}(M,L)=0$ as we have seen above. Thus if $\alpha^{c}(M,L)\neq 0$, there are two possibilities: either there is a pair $(g,A)$ with $\mu_{1}(\mathsf{L}_{(g,A)})=0$, or else $\mu_{1}(\mathsf{L}_{(g,A)})<0$ for any pair $(g,A)$.
The case when $\mu_{1}(\mathsf{L}_{(g,A)})=0$ (corresponding to the case (2)) is especially interesting: just as in the classic case, we prove a rigidity result as follows:
Theorem 1.5.
Let $(M,L)$ be a closed connected spinc manifold which is non-spin, and assume that $\alpha^{c}(M,L)\neq 0$. Suppose there are a metric $g$ on $M$ and a connection $A$ on $L$ for which $\mu_{1}(\mathsf{L}_{(g,A)})=0$. Then $(M,g)$ is pointwise conformal to another metric $g_{1}$ for which $R^{\text{{gen}}}_{(g_{1},A)}$ vanishes identically. Furthermore, $(M,L)$ with metric $g_{1}$ admits a parallel spinor, and $(M,g_{1})$ is a Riemannian product of a Kähler manifold and a simply connected spin manifold with a parallel spinor. In particular, if $M$ does not split as a product, then $M$ is conformally Kähler, and $L$ is either the canonical or the anti-canonical line bundle on $M$.
We also observe that there is a partial converse to Theorem 1.5 that applies to Kähler-Einstein Fano varieties. A very similar formulation appears in [11].
Theorem 1.6.
(cf. Goette-Semmelmann, [11, Theorem 1.9]) Let $(M,g)$ be a closed connected Kähler-Einstein manifold with positive Ricci curvature. Let $L$ be the anti-canonical line bundle on $M$. Then $(M,L)$ does not admit a connection $A$ on $L$ with $R^{\text{{gen}}}_{(g,A)}>0$, but with the connection $A$ on $L$ with harmonic curvature (necessarily a multiple of $g$), $R^{\text{{gen}}}_{(g,A)}\equiv 0$.
Next, we consider the space $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ of triples $(g,h,A)$ such that $R^{\text{{gen}}}_{(g,A)}>0$. Note that the space $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ coincides with the corresponding space $\mathcal{R}^{{\mathrm{tw}+}}(M,L)$ of triples $(g,h,A)$ such that $R^{\text{{tw}}}_{(g,A)}>0$.
In fact there is no loss of generality in fixing $h$ once and for all (see Proposition 3.2 below), but it would seem at first that it is important to consider variations not just in the Riemannian metric $g$, which determines the scalar curvature term $R_{g}$, but also in the connection $A$, which determines the extra curvature term $\mathcal{R}_{L}$. We will show, however, that this is not really necessary, and that once one has $(g,h,A)\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$, one can change the connection $A$ to a more canonical choice, only depending on the metric $g$, for which the positivity condition is still satisfied.
To study the topology of $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$, we will use a version of the “index difference” homomorphism,
$\operatorname{inddiff}(M)\colon\,\pi_{q}(\mathcal{R}^{+}(M))\to KO_{q+n+1}$
(where $M$ is a closed spin manifold, $\dim M=n\geq 5$, and $\mathcal{R}^{+}(M)$ is the space of psc metrics on $M$) studied in [7, 4, 9]. Loosely, it follows from [4] that the index difference homomorphism is non-trivial if the target group $KO_{q+n+1}$ is non-zero.
Similarly, if $(M,L)$ is a closed spinc manifold with a fixed choice of basepoint $(g_{0},A_{0})\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$, we define the spinc index difference homomorphism
(1.2) $\operatorname{inddiff}^{c}(M,L)\colon\,\pi_{q}(\mathcal{R}^{{\mathrm{gen}+}}(M,L))\to KU_{q+n+1}\,.$
We would like to address the following
Conjecture 1.7.
Let $(M,L)$ be a closed spinc non-spin manifold with $\dim M\geq 5$, and with $\mathcal{R}^{{\mathrm{gen}+}}(M,L)\neq\emptyset$. Then the index difference homomorphism (1.2) is non-trivial if the target group is non-zero.
Remark 1.8.
In fact there is an index-difference map $\mathcal{R}^{{\mathrm{gen}+}}(M,L)\to\Omega^{\infty+n}\mathbf{KU}$ (we use the notation from [7]). It is very likely that the techniques from [7, 4, 9] could be (and will be) generalized to the spinc case.
In this paper, we prove a couple of easy results providing evidence for Conjecture 1.7 to be true:
Theorem 1.9.
  1. (i)
    Let $(M,L)$ be a closed connected $n$-dimensional spinc manifold, with $n=4k-1\geq 7$, and $\mathcal{R}^{{\mathrm{gen}+}}(M,L)\neq\emptyset$. Then $\operatorname{inddiff}^{c}(M,L)\colon\,\pi_{0}(\mathcal{R}^{{\mathrm{gen}+}}(M,L))\to\mathbb{Z}$ is nonzero, and in fact, $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ has infinitely many connected components.
  2. (ii)
    There exists a closed simply connected spinc manifold $(M,L)$, not spin, for which infinitely many homotopy groups of $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ are non-zero.
1.2. Plan of the paper
In Section 2 we study the question: does there always exist a spinc line bundle $L$ with $\alpha^{c}(M,L)=0$? In Section 3 we analyze the relationship between the curvatures $R^{\text{{gen}}}$ and $R^{\text{{tw}}}$ as well as the behavior of $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ under conformal deformations. We prove the trichotomy theorem in Section 4 and analyze the rigidity case in Section 5. We define and study the index-difference map $\operatorname{inddiff}^{c}(M,L)$ in Section 6.
1.3. Acknowledgment
We would like to thank the referee for a careful reading of the paper and for many useful suggestions.
2. More About Spinc Manifolds
In this section we want to go into some more detail about spinc manifolds. First of all, we should clarify some terminology. Often the words “spinc manifold” are used to describe a manifold admitting a spinc structure, but really one should avoid this since spinc structures are rarely unique, and in fact never are in the case we are primarily interested in here, namely simply connected manifolds $M$ that are not spin but admit a spinc structure. That’s because $M$ being simply connected implies $H^{2}(M,\mathbb{Z})$ is torsion free, and $M$ being non-spin but admitting a spinc structure means $w_{2}(M)\neq 0$ and is the reduction of an integral class in $H^{2}(M,\mathbb{Z})$. Thus there must be an infinite set of cohomology classes, all representing distinct spinc structures, that reduce mod $2$ to $w_{2}(M)$. Each such cohomology class is the first Chern class $c_{1}(L)$ of a line bundle $L$, and it is best to refer to the pair $(M,L)$ as being a spinc manifold. (Of course $L$ is only determined up to isomorphism; isomorphism classes of line bundles, with the operation of tensor product of bundles, give the Picard group $\operatorname{Pic}(M)$, isomorphic to $H^{2}(M,\mathbb{Z})$ via the first Chern class $c_{1}$.)
Suppose $M^{n}$ is closed, simply connected, and non-spin, but admits a spinc structure $(M,L)$. Then the index $\alpha^{c}(M,L)$ of the spinc Dirac operator on $(M,L)$ is well-defined. Note that the index vanishes identically if $n$ is odd. Then the following question naturally arises:
Question 2.1.
Let $M$ be a closed simply connected manifold with even dimension $n=2k$. Fix an orientation on $M$. Assume that $w_{2}(M)\neq 0$ (so that $M$ is not spin) but that $w_{2}(M)$ is the reduction of an integral class (so that $M$ admits spinc structures $(M,L)$). Does there always exist a spinc line bundle $L$ with $\alpha^{c}(M,L)=0$? If so, how many such line bundles $L$ satisfy this condition?
The answer to this is the following:
Theorem 2.2.
Let $M$ be a simply connected closed manifold with even dimension $n=2k\geq 6$. Fix an orientation on $M$. Assume that $w_{2}(M)\neq 0$ but that $w_{2}(M)$ is the reduction of an integral class. If the second Betti number $\beta_{2}(M)$ is $1$ and there is a bundle $L$ with $\alpha^{c}(M,L)\neq 0$, then the number of bundles $L^{\prime}$ (up to isomorphism) with $\alpha^{c}(M,L^{\prime})=0$ is finite. Furthermore, there exist such manifolds $M$ with $\alpha^{c}(M,L)\neq 0$ for all spinc line bundles $L$. On other hand, there also exist such manifolds with $\alpha^{c}(M,L)=0$ for all spinc line bundles $L$.
Proof.
As we saw above, $H^{2}(M,\mathbb{Z})$ is torsion-free and infinite, and the even classes in this group (those reducing mod $2$ to 00) act simply transitively on the classes of spinc line bundles compatible with the orientation. Let $x=c_{1}(L)\in H^{2}(M,\mathbb{Z})$. By [19, Theorem D.15],
$\alpha^{c}(M,L)$ $=\left\langle\widehat{\mathbf{A}}(M)e^{x/2},[M]\right\rangle$
$=\left\langle\left(\frac{1}{k!}\left(\frac{x}{2}\right)^{k}+\cdots+1\right)\left(1-\frac{1}{24}p_{1}+\cdots\right),[M]\right\rangle$
$=\frac{1}{2^{k}\,k!}\langle x^{k},[M]\rangle+\text{lower order terms in }x,$
and so vanishing of $\alpha^{c}(M,L)$ is a polynomial condition on $x=c_{1}(L)$, which we can think of as living in the real vector space $H^{2}(M,\mathbb{R})$. If $\beta_{2}(M)=1$, this vector space is one-dimensional, and a polynomial equation in one variable can have only finitely many roots unless it vanishes identically. So this proves the first statement.
To get an example where $\alpha^{c}(M,L)=0$ for all $L$, consider $M=\mathbb{C}\mathbb{P}^{2}\times S^{4}$. If $x_{2}$ is the standard generator of $H^{2}(\mathbb{C}\mathbb{P}^{2},\mathbb{Z})$ and $y_{4}$ is the standard generator of $H^{4}(S^{4},\mathbb{Z})$, then $M$ has the cohomology ring $\mathbb{Z}[x_{2},y_{4}]/(x_{2}^{3},y_{4}^{2})$ and every class in $H^{2}(M,\mathbb{Z})$ is of the form $kx_{2}$. Furthermore, $e^{kx_{2}/2}=1+\frac{kx_{2}}{2}+\frac{k^{2}x_{2}^{2}}{8}$, while $\widehat{\mathbf{A}}(M)=1-\frac{x_{2}^{2}}{8}$. Since neither of these polynomials involves $y_{4}$, their product must evaluate to 00 on $[M]$.
Now we give an example where $M$ doesn’t admit any spinc line bundle $L$ with vanishing index. We take $M=\mathbb{C}\mathbb{P}^{4}\#N$, where $N$ is a $3$-connected spin $8$-manifold with $\widehat{A}(N)=1$. (One can construct such an $N$ via the $E_{8}$-plumbing; see [6].) Since $N$ is $3$-connected, any spinc line bundle on $M$ comes uniquely from one on $\mathbb{C}\mathbb{P}^{4}$. But for a line bundle $L$ on $\mathbb{C}\mathbb{P}^{4}$ with Chern class $c_{1}(L)=(2k+1)x$, where $x$ is the standard generator of $H^{2}(\mathbb{C}\mathbb{P}^{4},\mathbb{Z})$ — we write things this way since the Chern class must be odd by the spinc condition — we have
$\alpha^{c}(\mathbb{C}\mathbb{P}^{4},L)$ $=\langle\widehat{\mathbf{A}}(\mathbb{C}\mathbb{P}^{4})e^{(2k+1)x/2},[\mathbb{C}\mathbb{P}^{4}]\rangle$
$=\text{coefficient of }x^{4}\text{ in }\left(\frac{x/2}{\sinh(x/2)}\right)^{5}e^{(2k+1)x/2}$
$=\text{coefficient of }x^{4}\text{ in }\left(1-\frac{5x^{2}}{24}+\frac{3x^{4}}{128}\right)\times$
$\qquad\qquad\left(1+\frac{(2k+1)x}{2}+\frac{(2k+1)^{2}x^{2}}{8}+\frac{(2k+1)^{3}x^{3}}{48}+\frac{(2k+1)^{4}x^{4}}{384}\right)$
$=\frac{1}{24}(k-1)k(k+1)(k+2).$
(This calculation, by the way, is also in [12, Proposition 4.3] and in [11, Remark 3.2].) For $k$ an integer, this is always non-negative: it vanishes for $k=-2,-1,0,1$. We obtain
$\alpha^{c}(M,L)=\alpha^{c}(\mathbb{C}\mathbb{P}^{4},L)+\widehat{A}(N)\geq 1,$
which never vanishes. ∎
Remark 2.3.
If $\beta_{2}(M)>1$ and $\alpha^{c}(M,L)$ doesn’t vanish identically, then it can still happen that the number of line bundles $L$ with $\alpha^{c}(M,L)=0$ can be infinite. An example is $M=\mathbb{C}\mathbb{P}^{2}\#\overline{\mathbb{C}\mathbb{P}^{2}}$, $\mathbb{C}\mathbb{P}^{2}$ with one point blown up. We can identify $H^{2}(M,\mathbb{Z})$ with $\mathbb{Z}x\oplus\mathbb{Z}y$, where $x^{2}=-y^{2}$ is dual to $[M]$ and $xy=0$. Then if $c_{1}(L)=ax+by$, we have
$\alpha^{c}(M,L)$ $=\left\langle\widehat{\mathbf{A}}(M)e^{c_{1}(L)/2},[M]\right\rangle$
$=\left\langle\left(1-\frac{x^{2}}{8}+\frac{x^{2}}{8}\right)e^{ax/2}e^{by/2},[M]\right\rangle$
$=\frac{a^{2}-b^{2}}{8}.$
(Note that $a$ and $b$ both have to be odd to get a spinc structure.) This vanishes whenever $a=\pm b$, so there are infinitely many choices with vanishing $\alpha^{c}$.
3. Further Aspects of the “Generalized psc” Condition
Let $(M,L)$ be a spinc manifold, $h$ and $A$ a hermitian metric and unitary connection on the line bundle $L$, as usual. First we want to justify our previous assertion that there is no loss of generality in fixing the metric $h$ on $L$ once and for all.  [^3] We use this and hereafter omit $h$ from most notations. Before we do this, we need to fix some terminology. If $\omega$ is a (real) $2$-form on a manifold $M$, we denote by $|\omega|_{op}$ the pointwise operator norm of Clifford multiplication $c(\omega)$, acting on spinors. In other words, if $x\in M$,
$|\omega|_{op}(x)=\max\{|c(\omega)\psi(x)|:|\psi(x)|=1\},$
where the maximum is taken over all spinors $\psi$ with norm $1$ at $x$.
Lemma 3.1.
Let $(M,L)$ be a spinc manifold. Let $(g,h,A)$ be a triple ($g$ a Riemannian metric on $M$, $h$ a hermitian metric on $L$, and $A$ a connection on $L$). Then the condition $R^{\text{{tw}}}_{(g,A)}>0$ is equivalent to the condition $R^{\text{{gen}}}_{(g,A)}>0$.
Proof.
Clifford multiplication $c(\Omega)$ by $\Omega$ is skew-adjoint, and its eigenvalues come in pairs ±$\pm i\lambda$, with $\lambda$ real, so the condition that $R^{\text{{tw}}}_{(g,A)}=R_{g}+2\mathcal{R}_{L}>0$ reduces to saying that $R_{g}>2|\Omega|_{op}$, where $|\Omega|_{op}$ is the pointwise operator norm of $c(\Omega)$, or that the eigenvalue $i\lambda$ of $c(\Omega)$ at any point $x\in M$, with $\lambda\geq 0$ as big as possible, is less than $\frac{1}{2}R_{g}(x)$. This is the condition $R^{\text{{gen}}}_{(g,A)}>0$.
In the other direction, if $R^{\text{{gen}}}_{(g,A)}>0$, then the minimal eigenvalue of $ic(\Omega)$ is everywhere less than $\frac{1}{2}R_{g}(x)$ in absolute value, which shows that the self-adjoint operator $R^{\text{{tw}}}_{(g,A)}$ is positive. ∎
Thus Lemma 3.1 says that $R^{\text{{tw}}}_{(g,A)}>0\Leftrightarrow R^{\text{{gen}}}_{(g,A)}>0$, with the latter a more manageable condition to work with since $R^{\text{{gen}}}_{(g,A)}$ is scalar-valued. Note that both conditions imply vanishing of $\alpha^{c}(M,L)$, because of equation (1.1). We denote by $\mathcal{R}(M)$ the space of Riemannian metrics on $M$, by $\mathcal{A}(M,L)$ the space of connections on $L$ and by $\mathcal{H}(L)$ the space of Hermitian metrics on $L$. Let
$\mathcal{R}^{{\mathrm{tw}+}}(M,L):=\bigl{\{}(g,A,h)\in\mathcal{R}(M)\times\mathcal{A}(M,L)\times\mathcal{H}(L)\mid R^{\text{{tw}}}_{(g,A,h)}>0\bigr{\}}.$
Let $h_{0}\in\mathcal{H}(L)$ be a fixed Hermitian metric on $L$, and let $\mathcal{R}^{{\mathrm{tw}+}}(M,L)_{h_{0}}=\{(g,A,h_{0})\in\mathcal{R}^{{\mathrm{tw}+}}(M,L)\}$. Since the contractible group $C^{\infty}_{+}(M)$ of positive smooth functions acts freely and transitively on $\mathcal{H}(L)$ by pointwise multiplication, for any metric $h\in\mathcal{H}(L)$, there is a unique function $f\in C^{\infty}_{+}(M)$ such that $f\cdot h_{0}=h$. We claim that if $R^{\text{{tw}}}_{(g,A,h_{0})}>0$ for some triple $(g,A,h_{0})$, then $R^{\text{{tw}}}_{(g,A,f\cdot h_{0})}>0$ for any $f\in C^{\infty}_{+}(M)$. To see this, observe that changing the metric does not change the curvature $2$-form $\Omega$ or the Clifford algebra bundle, though it does change the pointwise norm on spinors. At the same time, the positivity conditions of Lemma 3.1 do not change when the norm on $L$ is rescaled by the constant $f(x)$. Thus we have proved the following:
Proposition 3.2.
Let $(M,L)$ be a spinc manifold with $\mathcal{R}^{{\mathrm{tw}+}}(M,L)\neq\emptyset$, and let $h_{0}\in\mathcal{H}(L)$ be a fixed Hermitian metric. Then the inclusion $\mathcal{R}^{{\mathrm{tw}+}}(M,L)_{h_{0}}\subset\mathcal{R}^{{\mathrm{tw}+}}(M,L)$ is a homotopy equivalence.
Remark 3.3.
From now on, we identify the spaces $\mathcal{R}^{{\mathrm{tw}+}}(M,L)$ and $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$, and because of Proposition 3.2 we will work with the space $\mathcal{R}^{{\mathrm{tw}+}}(M,L)$ assuming that the line bundle $L$ and the metric $h$ on $L$ are fixed once and for all. We will omit the metric $h$ from the notations, i.e., work with pairs $(g,A)$ as elements of $\mathcal{R}^{{\mathrm{tw}+}}(M,L)$ and $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$. Furthermore, we will see soon that we really only need to keep track of the Riemannian metric $g$ and not of the bundle connection $A$ since there is a canonical choice of the connection which makes the curvature 2-form $\Omega$ harmonic with respect to the metric $g$.
Next, we want to somewhat simplify the study of pairs $(g,A)\in\mathcal{R}^{{\mathrm{tw}+}}(M,L)$. As a warm-up, let’s consider just the usual psc condition. Let $M$ be a closed $n$-manifold admitting a psc metric, $n\geq 3$. Let $\mathcal{R}(M)$ be the space of all Riemannian metrics on $M$, and let $\mathcal{R}^{+}(M)\subset\mathcal{R}(M)$ be the subspace of psc metrics. The space of (pointwise) conformal classes $\mathcal{C}(M)$ is defined as the orbit space of the left action
$C^{\infty}_{+}(M)\times\mathcal{R}(M)\to\mathcal{R}(M),\ \ \ \ (f,g)\mapsto f\cdot g,$
so that $\mathcal{C}(M):=C^{\infty}_{+}(M)\backslash\mathcal{R}(M)$ (here $C^{\infty}_{+}(M)$ is the space of positive smooth functions). Denote by $p\colon\,\mathcal{R}(M)\to\mathcal{C}(M)$ the corresponding projection. We give $\mathcal{C}(M)$ the quotient topology. Note that $\mathcal{C}(M)$ is Hausdorff. To see this, it suffices to show that if $g_{n}\in\mathcal{R}(M)$, $f_{n}\in C^{\infty}_{+}(M)$, and $g_{n}\to g_{0}$, $f_{n}g_{n}\to g_{1}$, then $g_{1}=fg_{0}$ with $f_{n}\to f$ for some $f$ in $C^{\infty}_{+}(M)$. But this is clearly true pointwise, and since $\{g_{n}\}$ and $\{f_{n}g_{n}\}$ converge uniformly along with their derivatives, by definition of the topology on $\mathcal{R}(M)$, it is also clear that one gets such convergence of $f_{n}$ to $f$. We say that a conformal class $C\in\mathcal{C}(M)$ is positive if it contains a psc-metric, and let $\mathcal{C}^{+}(M)\subset\mathcal{C}(M)$ be the subspace of positive conformal classes. We note that the subspace $\mathcal{R}^{+,c}(M):=p^{-1}(\mathcal{C}^{+}(M))$ of “conformally positive Riemannian metrics” $g$ is strictly larger then the space $\mathcal{R}^{+}(M)$.
According to Kazdan-Warner [18, Theorem 3.2], the sign of the first eigenvalue $\mu_{1}(\mathsf{L}_{g})$ of the conformal Laplacian
(3.1) $\mathsf{L}_{g}=-\frac{4(n-1)}{n-2}\Delta_{g}+R_{g}$
is constant on conformal classes, and the space $\mathcal{R}^{+,c}(M)$ can be characterized as the set of metrics for which $\mu_{1}(\mathsf{L}_{g})$ is strictly positive. Indeed, if $\mu_{1}(\mathsf{L}_{g})>0$, then the corresponding eigenfunction $\varphi$ can be chosen strictly positive and [18, Remark 2.12] the (pointwise) conformal metric $g_{1}=\varphi^{\frac{4}{n-2}}g$ has $R_{g_{1}}>0$ everywhere. Conversely, if $R_{g}>0$, then it is clear that $\mathsf{L}_{g}$ is bounded away from zero by a positive constant (the minimum value of $R_{g}$), and thus $\mu_{1}(\mathsf{L}_{g})>0$. So $\mathcal{R}^{+}(M)\subseteq\mathcal{R}^{+,c}(M)$ and $\mathcal{R}^{+,c}(M)$ is the smallest family of Riemannian metrics containing $\mathcal{R}^{+}(M)$ and closed under (pointwise) conformal changes of the metric. On spin manifolds, invertibility of the Dirac operator extends from $\mathcal{R}^{+}(M)$ to $\mathcal{R}^{+,c}(M)$ via [14], which says that $\lambda_{1}^{2}(D_{g})\geq\frac{n}{4(n-1)}\mu_{1}(\mathsf{L}_{g})>0$, where $\lambda_{1}(D_{g})$ is the first eigenvalue of the Dirac operator $D_{g}$.
Theorem 3.4.
Let $M$ be a closed manifold, $\dim M=n\geq 3$, admitting a metric of positive scalar curvature. The inclusion $\mathcal{R}^{+}(M)\hookrightarrow\mathcal{R}^{+,c}(M)$ is a homotopy equivalence.
Proof.
Note that $\mathcal{R}^{+}(M)$ and $\mathcal{R}^{+,c}(M)$ are both open subsets of the separable Fréchet space of symmetric $2$-tensors on the tangent bundle, so they each have the homotopy type of a countable CW complex [20, Theorem 1]. We will show that the projection maps $\mathcal{R}^{+}(M)\to{\mathcal{C}^{+}(M)}$ and $p\colon\,\mathcal{R}^{+,c}(M)\to{\mathcal{C}^{+}(M)}$ are both fibrations with contractible fibers, and hence are homotopy equivalences. By the definitions, both of these maps are surjective, and $\mathcal{C}^{+}(M)$ has the quotient topology. Since $C^{\infty}_{+}(M)$ is a contractible Fréchet Lie group (under multiplication) and it acts smoothly and freely on $\mathcal{R}^{+,c}(M)$, with $\mathcal{C}^{+}(M)$ the (paracompact) quotient space, the projection map $p\colon\,\mathcal{R}^{+,c}(M)\to\mathcal{C}^{+}(M)$ is a principal bundle for a contractible Fréchet Lie group, and thus is a homotopy equivalence. In particular, it has a section. Choose such a section $s_{1}\colon\,\mathcal{C}^{+}(M)\to\mathcal{R}^{+,c}(M)$; it’s not canonical, but that won’t matter for us. So for each conformal class $C\in\mathcal{C}^{+}(M)$, $s_{1}(C)$ is a metric in the class $C$ for which $\mu_{1}(\mathsf{L}_{s_{1}(C)})>0$. There is a positive eigenfunction $\varphi_{C}$ for the eigenvalue $\mu_{1}(\mathsf{L}_{s_{1}(C)})$, and it’s unique up to multiplication by positive scalars. We can make it unique (on the nose!) by normalizing it to have maximum value $1$ on $M$. Then the map $C\mapsto\varphi_{C}$ is continuous in the appropriate topologies, because of elliptic regularity. Now we get a section $s\colon\,\mathcal{C}^{+}(M)\to\mathcal{R}^{+}(M)$ by taking $s(C)=\varphi_{C}^{\frac{4}{n-2}}s_{1}(C)$, using [18, Remark 2.12] to confirm that the scalar curvature of this metric is indeed strictly positive. We need to show that $\mathcal{R}^{+}(M)\to\mathcal{C}^{+}(M)$ is also a fibration with contractible fibers. We will show not only that the fibers are contractible open subsets of the fibers of $p$, but also that there is a deformation retraction from $\mathcal{R}^{+}(M)$ down to $s(\mathcal{C}^{+}(M))$ (a homeomorphic copy of $\mathcal{C}^{+}(M)$), and thus that $\mathcal{R}^{+}(M)$ and $\mathcal{C}^{+}(M)$ are homotopy equivalent.
Fix $C\in\mathcal{C}^{+}(M)$ and some metric $g_{0}\in p^{-1}(C)\cap\mathcal{R}^{+}(M)$. We will show that $p^{-1}(C)\cap\mathcal{R}^{+}(M)$ is convex in a suitable sense, hence also contractible.
[…]
Now we get a deformation retraction of $\mathcal{R}^{+}(M)$ down to $s(\mathcal{C}^{+}(M))$ by using the above homotopy from a general metric $g_{1}\in\mathcal{R}^{+}(M)$ to $g_{0}=s(p(g_{1}))$, and thus $p$ restricted to $\mathcal{R}^{+}(M)$ is a homotopy equivalence. ∎
The advantage of Theorem 3.4 is that it makes it possible to study the topology of $\mathcal{R}^{+}(M)$ via tools involving conformal geometry.
Now we would like to do something similar for the curvature condition $R^{\text{{gen}}}_{(g,A)}>0$ on spinc manifolds (when $n\geq 3$). For this the paper [14], giving estimates on eigenvalues of the Dirac operator using the conformal Laplacian, should be replaced by [3, 13, 22]. What Bär calls the curvature endomorphism $\mathcal{K}$ is $\frac{1}{4}R_{g}+\frac{1}{2}ic(\Omega)=\frac{1}{4}R^{\text{{tw}}}_{(g,A)}$, which is a self-adjoint endomorphism at each point of $M$, and he takes $\kappa$ [3, bottom of p. 42] to be a scalar lower bound for $\mathcal{K}$, which since the eigenvalues of $ic(\Omega)$ come in $\pm$ pairs can be taken to be $\frac{1}{4}R^{\text{{gen}}}_{(g,A)}$, so the estimate from [3, Theorem 3] turns out to be exactly the inequality
$\lambda_{1}^{2}(D_{L})\geq\frac{n}{4(n-1)}\mu_{1}(\mathsf{L}_{(g,A)}),$
with $\mathsf{L}_{(g,A)}$ defined by replacing $R_{g}$ by $R^{\text{{gen}}}_{(g,A)}$ in the formula for $\mathsf{L}_{g}$, and this is nothing but the spinc analogue of Hijazi’s inequality.
In some cases, one can basically get rid of the connection $A$ on the line bundle $L$ for the condition $R^{\text{{gen}}}_{(g,A)}>0$. Here is a case where this is possible.
Lemma 3.5.
Let $L$ be a line bundle on a closed Riemannian manifold $(M,g)$. Then the $2$-norm [^4] This is defined by $\|\Omega\|_{2}^{2}=\int_{M}\Omega\wedge*\Omega$. of the curvature $2$-form $\Omega$ of a connection on $L$ is minimized when $\Omega$ is the unique harmonic $2$-form in its de Rham class.
Proof.
This fact is well-known but we recall the proof. Recall the de Rham class of the form $\Omega$ (representing the first Chern class $c_{1}(L)=[\frac{1}{2\pi}\Omega]$ in $H^{2}(M,\mathbb{R})$) contains a unique harmonic $2$-form $\Omega_{0}$, which (up to a factor of $2\pi$) is the curvature of a connection $A_{0}$ on $L$. Any other connection on $L$ is of the form $A_{0}+\alpha$, where $\alpha$ is an ordinary $1$-form, and has curvature $2$-form $\Omega=\Omega_{0}+d\alpha$. We have for any $1$-form $\alpha$
$\langle\Omega_{0},d\alpha\rangle=\langle d^{*}\Omega_{0},\alpha\rangle=0$
since $d^{*}\Omega_{0}=0$ because $\Omega_{0}$ is harmonic. Thus, by the geometry of inner product spaces,
$\|\Omega\|_{2}^{2}=\|\Omega_{0}\|_{2}^{2}+\|d\alpha\|_{2}^{2}$
is minimized when $d\alpha=0$ and $\Omega=\Omega_{0}$ is harmonic. (See also [19, Lemma D.13].) ∎
Remark 3.6.
Clearly the condition $R^{\text{{gen}}}_{(g,A)}=R_{g}-2|\Omega|_{op}>0$ is not directly a condition on the $2$-norm of $\Omega$. However, by [13, Lemma 3.3] and the fact that Clifford multiplication by $\Omega$ is skew-symmetric, one has the pointwise estimate $|\Omega|_{op}\leq\sqrt{\frac{n}{2}}|\Omega|_{2}$. In the other direction, one has a similar estimate
$|\Omega|_{2}^{2}=\frac{1}{2}\operatorname{Tr}c(\Omega)^{*}c(\Omega)\leq\frac{n}{2}|\Omega|_{op}^{2}$
and thus $|\Omega|_{2}\leq\sqrt{\frac{n}{2}}|\Omega|_{op}$. If $\Omega$ is harmonic, then $\Omega$ has constant coefficients in a local harmonic coordinate system, and thus $|\Omega|_{2}$ and $|\Omega|_{op}$ are constant functions. Furthermore, one always has
[…]
which when $\Omega$ is harmonic becomes $\text{{vol}}(M)\,|\Omega|_{2}^{2}$. Thus in this case $\|\Omega\|_{2}=\text{{vol}}(M)^{\frac{1}{2}}|\Omega|_{2}$, and if $\|\Omega\|_{2}$ is small enough, one can use Lemma 3.5 to replace the original generalized psc connection by one with harmonic curvature. But the estimates comparing $|\Omega|_{op}$ and $|\Omega|_{2}$ get worse and worse as $n\to\infty$.
Now we proceed to try to find an analogue of Theorem 3.4. Let $(M,L)$ be a closed spinc manifold of dimension $n\geq 3$ and $\mathcal{R}^{{\mathrm{gen}+}}(M,L)\neq\emptyset$. Let $\Omega$ be the curvature $2$-form on $A$. Define the modified conformal Laplacian of $(M,L)$ equipped with a Riemannian metric $g$ and a connection $A$ to be
(3.2) $\mathsf{L}_{(g,A)}:=-\frac{4(n-1)}{n-2}\Delta_{g}+R^{\text{{gen}}}_{(g,A)}=\mathsf{L}_{g}-2|\Omega|_{op}.$
Note of course that if $M$ is spin, the line bundle $L$ is trivial and we can take $\Omega=0$, so then $R^{\text{{gen}}}_{(g,A)}=R_{g}$ and $\mathsf{L}_{(g,A)}$ is just the usual conformal Laplacian $\mathsf{L}_{g}$. By [3, Theorem 3] (even in the general spinc case), the lowest eigenvalue $\mu_{1}(\mathsf{L}_{(g,A)})$ of $\mathsf{L}_{(g,A)}$ gives a lower bound for the smallest eigenvalue of the square of the spinc Dirac operator $D_{L}$:
(3.3) $\lambda_{1}^{2}(D_{L})\geq\frac{n}{4(n-1)}\mu_{1}(\mathsf{L}_{(g,A)}).$
Let’s show that $R^{\text{{gen}}}_{(g,A)}\geq 0$ with $R^{\text{{gen}}}_{(g,A)}$ positive somewhere on $M$ implies that $\mu_{1}(\mathsf{L}_{(g,A)})>0$. The argument is the same as for $\mathsf{L}_{g}$. Suppose $R^{\text{{gen}}}_{(g,A)}\geq 0$ and $R^{\text{{gen}}}_{(g,A)}$ is positive somewhere. Clearly $\mathsf{L}_{(g,A)}\geq 0$, so that $\mu_{1}(\mathsf{L}_{(g,A)})\geq 0$. But suppose $\mu_{1}(\mathsf{L}_{(g,A)})=0$. There is always positive eigenfunction for $\mu_{1}(\mathsf{L}_{(g,A)})$, so we have a strictly positive function $\phi$ in the kernel of $\mathsf{L}_{(g,A)}$. We have
[…]
and both terms on the right are $\geq 0$, so they both vanish. Vanishing of the first term shows that $\phi$ is a positive constant, and then vanishing of the second term shows that $R^{\text{{gen}}}_{(g,A)}\equiv 0$, a contradiction. In particular, the inequality (3.3) implies that the Dirac operator $D_{L}$ is invertible.
We denote by $\mathcal{R}^{{\mathrm{gen}+},c}(M,L)$ the space of pairs $(g,A)\in\mathcal{R}(M)\times\mathcal{A}(M,L)$ such that $\mu_{1}(\mathsf{L}_{(g,A)})>0$. Now we consider what happens under pointwise conformal deformation of the metric $g$. Note that we don’t change $h$ or $A$, so $\Omega$ is fixed, though the norm $|\Omega|_{op}$ of $\Omega$ varies with metric, since it depends on the Clifford algebra.
Lemma 3.7.
Let $(M,L)$ be a closed spinc manifold, $\dim M=n\geq 3$, and let $(g,A)$ be a pair such that $\mu_{1}(\mathsf{L}_{(g,A)})>0$. Then there exists a conformal metric $g_{1}=u^{\frac{4}{n-2}}g$ such that $R^{\text{{gen}}}_{(g_{1},A)}>0$. Moreover, the space $\mathcal{R}^{{\mathrm{gen}+},c}(M,L)$ is invariant under pointwise conformal changes of the metric, and for $(g,A)\in\mathcal{R}^{{\mathrm{gen}+},c}(M,L)$, there is a pointwise conformal metric $g_{1}$ with $R^{\text{{gen}}}_{(g_{1},A)}>0$
Proof.
Let $u\in C^{\infty}(M)$, $u>0$, and let $g_{1}=u^{\frac{4}{n-2}}g$. We need to show that $R^{\text{{gen}}}_{(g_{1},A)}>0$. We have $R_{g_{1}}=u^{-\alpha}\mathsf{L}_{g}(u)$, where $\alpha=1+\frac{4}{n-2}$. If we rescale the metric by a constant $c>0$, the curvature is multiplied by $c^{-1}$, and $|\Omega|_{op}$ scales the same way as the curvature, so that computed with respect to the new metric $g_{1}$, $|\Omega|_{op}$ is multiplied by $u^{-\frac{4}{n-2}}$. Then we use (3.2) to compute $R^{\text{{gen}}}_{(g_{1},A)}$:
$R^{\text{{gen}}}_{(g_{1},A)}=R_{g_{1}}-2|\Omega|_{op,g_{1}}$ $=u^{-\alpha}\mathsf{L}_{g}(u)-2u^{-\frac{4}{n-2}}|\Omega|_{op,g}$
$=-\frac{4(n-1)}{n-2}u^{-\alpha}\Delta_{g}(u)+u^{-\alpha}R_{g}u-2uu^{-\alpha}|\Omega|_{op,g}$
$=u^{-\alpha}{\mathsf{L}_{(g,A)}}(u).$
Finally, suppose $(g,A)\in\mathcal{R}^{{\mathrm{gen}+},c}(M,L)$, and let $\varphi$ be a strictly positive eigenfunction for the first eigenvalue $\mu_{1}(\mathsf{L}_{(g,A)})>0$. Replace $g$ by the pointwise conformal metric $g_{1}=\varphi^{\frac{4}{n-2}}g$. Then
$R^{\text{{gen}}}_{(g_{1},A)}=R_{g_{1}}-2|\Omega|_{op,g_{1}}$ $=\varphi^{-\alpha}\mathsf{L}_{g}(\varphi)-2\varphi^{-\frac{4}{n-2}}|\Omega|_{op,g}$
$=-\frac{4(n-1)}{n-2}\varphi^{-\alpha}\Delta_{g}(\varphi)+\varphi^{-\alpha}R_{g}\varphi-2\varphi\varphi^{-\alpha}|\Omega|_{op,g}$
$=\varphi^{-\alpha}\mathsf{L}_{(g,A)}(\varphi)=\mu_{1}(\mathsf{L}_{(g,A)})\varphi\varphi^{-\alpha}>0.$
This concludes the proof. ∎
Now we have an analogue of Theorem 3.4.
Theorem 3.8.
Let $(M,L)$ be a closed spinc manifold, $n\geq 3$, such that $\mathcal{R}^{{\mathrm{gen}+},c}(M,L)\neq\emptyset$. Then the inclusion $\mathcal{R}^{{\mathrm{gen}+}}(M,L)\subset\mathcal{R}^{{\mathrm{gen}+},c}(M,L)$ is homotopy equivalence.
Proof.
We proceed as in the proof of Theorem 3.4. Let $(g,A)\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$, i.e., $R^{\text{{gen}}}_{(g,A)}=R_{g}-2|\Omega|_{op}>0$, and let $g_{1}=u^{\frac{4}{n-2}}g_{0}$ be pointwise conformal to $g$, for some $u\in C^{\infty}_{+}(M)$. We have a homotopy $u_{t}=(1-t)+tu$, $g_{t}=u_{t}^{\frac{4}{n-2}}g_{0}$, $0\leq t\leq 1$, where $u_{t}\in C^{\infty}_{+}(M)$. By the formula for scalar curvature of a conformal deformation of a metric, $R_{g_{t}}=u_{t}^{-\alpha}\mathsf{L}_{g_{0}}(u_{t})$, where $\alpha=\frac{n+2}{n-2}$. But by linearity of $\mathsf{L}_{g_{0}}$, $\mathsf{L}_{g_{0}}(u_{t})=(1-t)\mathsf{L}_{g_{0}}(1)+t\mathsf{L}_{g_{0}}(u)=(1-t)R_{g}+t\mathsf{L}_{g_{0}}(u)$. So
$R^{\text{{gen}}}_{(g_{t},A)}=R_{g_{t}}-2|\Omega|_{op,g_{t}}$ $=u_{t}^{-\alpha}(1-t)R_{g}+u_{t}^{-\alpha}t\mathsf{L}_{g_{0}}(u)-2|\Omega|_{op,g}u_{t}u_{t}^{-\alpha}$
$=u_{t}^{-\alpha}\bigl{(}(1-t)R_{g}+t\mathsf{L}_{g_{0}}(u)-2|\Omega|_{op,g}(1-t)-2|\Omega|_{op,g}tu\bigr{)}$
$=u_{t}^{-\alpha}\bigl{(}(1-t)R^{\text{{gen}}}_{(g,A)}+t\mathsf{L}_{(g_{0},A)}(u)\bigr{)}>0,$
where $\kappa=R^{\text{{gen}}}_{(g_{0},A)}$. This shows that the fibers of the projection to pointwise conformal classes are convex, and so the projection map is a homotopy equivalence, by exactly the same argument as in the proof of Theorem 3.4. ∎
Now we can effectively get rid of having to worry about the connection on the line bundle by the following trick. Start with a pair $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}>0$. By Theorem 3.8 and the solution of the Yamabe problem, we can make a pointwise conformal change of the metric $g$ to a Riemannian metric $\check{g}$ with constant positive scalar curvature, and still stay in the space $\mathcal{R}^{{\mathrm{gen}+},c}(M,L)$. Now replace $g$ by $\check{g}$. The curvature $2$-form $\Omega$ must have the form $\Omega_{0}+d\alpha$, where $\Omega_{0}$ is the unique harmonic $2$-form (with respect to the metric $\check{g}$) in the de Rham class of $\Omega$. In harmonic coordinates (with respect to the metric $\check{g}$), $\Omega_{0}$ is a $2$-form with constant coefficients, so its pointwise norm $|\Omega_{0}|_{op,\check{g}}$ is a constant function. Thus if $|\Omega|_{op,\check{g}}$ is not constant, replacing $\Omega$ by $\Omega_{0}$ (which corresponds to replacing the initial connection $A$ by a connection $A_{0}$) lowers the maximum value of the pointwise operator norm, and replaces $R^{\text{{gen}}}_{(\check{g},A)}$ by a constant $R^{\text{{gen}}}_{(\check{g},A_{0})}=R_{\check{g}}-2|\Omega_{0}|_{op,\check{g}}$, thereby increasing its minimum value. Since $\mathsf{L}_{(\check{g},A_{0})}$ is positive, this constant must be positive.
Thus it is no loss of generality to assume that the curvature $2$-form is harmonic, and thus determined by its Chern class and the Riemannian metric.
Remark 3.9.
We note that the aforementioned argument is not applicable to an arbitrary family of initial metrics, $g_{\upsilon}$, $\upsilon\in\Upsilon$, because the Yamabe problem generally does not admit a unique solution.
4. A Trichotomy Theorem
In the last section, we saw that there is a close analogy between generalized scalar curvature on a spinc manifold and scalar curvature in the spin case. We state again a trichotomy theorem for spinc manifolds. Our corresponding theorem is the following:
Theorem 4.1.
Every closed connected spinc manifold $(M,L)$ of dimension at least three falls into exactly one of the following three classes:
  1. (1)
    those admitting a pair $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}>0$, in which case every smooth function on $M$ is the generalized scalar curvature of some pair $(g^{\prime},A^{\prime})$, $g^{\prime}$ a Riemannian metric on $M$ and $A^{\prime}$ a connection on $L$;
  2. (2)
    those admitting a pair $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}\geq 0$ but not a pair $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}$ positive everywhere — in this case a smooth function on $M$ is the generalized scalar curvature of some pair $(g^{\prime},A^{\prime})$ if and only if it’s either identically 00 or else negative somewhere;
  3. (3)
    all other manifolds, those not admitting any $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}\geq 0$ — in this case, a smooth function on $M$ is the generalized scalar curvature of some pair $(g^{\prime},A^{\prime})$ if and only if it is negative somewhere on $M$.
Note that in all three cases, any smooth function which is negative somewhere can be a generalized scalar curvature $R^{\text{{gen}}}_{(g,A)}$ for some $(g,A)$. This is Theorem 4.2 below. Rigidity aspects of case (2) of Theorem 4.1 will be discussed further in section 5. The proof of Theorem 4.1 will be broken up in several steps. The first step, helping to distinguish between cases (1) and (2), was already mentioned in section 3: if $R^{\text{{gen}}}_{g,A}\geq 0$ and is not identically 00, then $\mu_{1}(\mathsf{L}_{(g,A)})>0$, and so by pointwise conformal deformation of the metric from $g$ to $g^{\prime}$, one can arrange to make $R^{\text{{gen}}}_{(g^{\prime},A)}>0$ everywhere. The next step is the following, which clearly takes care of statements (2) and (3) in Theorem 4.1. The argument here is closely modeled on [18].
Theorem 4.2.
Let $(M,L)$ be a connected closed spinc manifold of dimension $n\geq 3$ and let $\kappa$ be a smooth function on $M$ taking a negative value somewhere on $M$. Then $\kappa$ is the generalized scalar curvature of some pair $(g,A)$, $g$ a Riemannian metric on $M$ and $A$ a connection on $L$.
Proof.
By [18, Lemma 3.6 and Theorem 4.1] or [15, Lemma 3(a)], $M$ always has a metric $g_{0}$ with constant negative scalar curvature. Choose any connection $A$ on $L$. Then $R^{\text{{gen}}}_{(g_{0},A)}$ is strictly negative everywhere. Consider the twisted conformal Laplacian (3.2) $\mathsf{L}_{(g_{0},A)}$. The lowest eigenvalue $\mu_{1}(\mathsf{L}_{(g_{0},A)})$ of this operator must be negative, since we have
[…]
(In fact the same calculation shows that
[…]
for any $(g,A)$ will imply that $\mu_{1}(\mathsf{L}_{(g_{0},A)})<0$.)
In order to make the arguments below transparent, let us briefly recall a relevant result by Kazdan-Warner from [18, Section 2]. They consider a general equation
(4.1) ${-\mathbf{L}u=\Delta u-hu=-Hu^{\alpha}\ ,}$
where $h$ and $H$ are prescribed functions and $\alpha>1$ is a constant. In our case we assume $h$ and $H$ are both smooth. We denote by $\mu_{1}(\mathbf{L})$ the first eigenvalue of $\mathbf{L}$, and by $u_{1}>0$ the corresponding eigenfunction. One can observe that if there exists a solution $u>0$ of (4.1), then the signs of $\mu_{1}(\mathbf{L})$ and $H$ are the same, provided $H$ is never zero; see [18, Lemma 2.5]. Then the following result holds:
Proposition 4.3.
[18, Theorem 2.11(b)] Assume that $H<0$; then a positive solution of (4.1) exists if and only if $\mu_{1}(\mathbf{L})<0$.
The idea to construct the solution $u$ is to use upper and lower solutions $u_{+}$ and $u_{-}$. In the case at hand, $u_{+}=$ large constant, and $u_{-}=\epsilon u_{1}$, where $u_{1}$ is the first eigenfunction: $\mathbf{L}u_{1}=\mu_{1}(\mathbf{L})u_{1}$.
Now we continue with the proof of Theorem 4.2. First suppose that the function $\kappa$ is everywhere negative. By [18, Theorem 2.11(b)] (or by Proposition 4.3), there is a smooth function $u>0$ such that with $g=u^{\frac{4}{n-2}}g_{0}$, $(g,A)$ has $R^{\text{{gen}}}_{(g,A)}=\kappa$, since by the calculation in the proof of Lemma 3.7, if we let $\alpha=\frac{n+2}{n-2}$, then we need to solve the equation $L_{(g_{0},A)}(u)=\kappa u^{\alpha}$, which is the same as
$R^{\text{{gen}}}_{(g,A)}=u^{-\alpha}L_{(g_{0},A)}(u)=\kappa.$
Then the existence of the solution $u$ follows from Proposition 4.3.
The case where $\kappa$ is allowed to be non-negative on some subset of $M$ is more delicate, since a pointwise conformal deformation of $g_{0}$ may not suffice. In this case we show first that after moving $\kappa$ around by a diffeomorphism, we can make it very negative on a “large” portion of $M$.
Let $g_{0}$ be the same metric on $M$ with constant negative scalar curvature. We choose a diffeomorphism $\phi\colon\,M\to M$ such that
[…]
This will suffice since if $R^{\text{{gen}}}_{(g,A)}=\phi^{*}(\kappa)$ for some diffeomorphism $\phi$ of $M$, then
$R^{\text{{gen}}}_{\bigl{(}(\phi^{-1})^{*}g,(\phi^{-1})^{*}(A)\bigr{)}}=\kappa.$
It is not so difficult to arrange a diffeomorphism $\phi$ such that the inequality (4.2) holds. Choose a constant $C_{-}>0$ such that $\kappa(x)<-C_{-}$ for some $x\in M$, and find a diffeomorphism $\phi$ such that the image $\phi(\{\kappa<-C_{-}\})$ has almost the same $g_{0}$-volume as the $g_{0}$-volume of $M$. If, say,
$\mathrm{Vol}_{g_{0}}(M)-\mathrm{Vol}_{g_{0}}(\phi(\{\kappa<-C_{-}\}))=\epsilon$
and $\kappa\leq C_{+}$, then
[…]
for small enough $\epsilon>0$. We note that the condition (4.2) means that the function $\phi^{*}(\kappa)$ is “not too positive too often”, see [18, p. 124 and Lemma 2.15]. This fact, combined with the condition that $\mu_{1}(\mathsf{L}_{(g_{0},A)})<0$, guarantees the existence of a smooth, positive function $u$ satisfying the equation $\mathsf{L}_{(g_{0},A)}u=\phi^{*}(\kappa)u^{\alpha}$. Thus we obtain $R^{\text{{gen}}}_{(g,A)}=\phi^{*}(\kappa)$ for suitable $\kappa$ and $u>0$, with $g=u^{\frac{4}{n-2}}g_{0}$, follows by the same continuity argument as in [18, §4]. ∎
End of the proof of Theorem 4.1.
To finish the proof of Theorem 4.1, we need to show that if $\kappa\geq 0$ and $(M,L)$ admits a pair $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}>0$, then $\kappa$ can be realized as the generalized scalar curvature of some pair $(g^{\prime},A^{\prime})$ on $(M,L)$. (The case where $\kappa$ is negative somewhere is already covered by Theorem 4.2.)
By the argument at the end of section 3, the assumption implies that $(M,L)$ admits a pair $(\check{g},A)$ with a constant $R_{\check{g}}$, and there is a connection $A$ on $M$ with curvature form $\Omega$ with $|\Omega|_{op,\check{g}}$ constant so that $R^{\text{{gen}}}_{(\check{g},A)}$ is a positive constant. By rescaling, we can obviously arrange for any positive constant to be realized as the function $R^{\text{{gen}}}_{(g^{\prime},A^{\prime})}$ for some pair $(g^{\prime},A^{\prime})$. To show that one can achieve $R^{\text{{gen}}}_{(g^{\prime},A^{\prime})}\equiv 0$ for some pair $(g^{\prime},A^{\prime})$, just argue as in the proofs of [16, Theorems 3.9 and 1.3]. So we just need to deal with the case where $\kappa\geq 0$ is not constant. Choose $(g^{\prime},A^{\prime})$ as just explained so that $R^{\text{{gen}}}_{(g^{\prime},A^{\prime})}$ is a positive constant. Since $\kappa$ is assumed $\geq 0$ and nonconstant, we can choose a constant $c>0$ so that $0\leq\min\kappa<cR^{\text{{gen}}}_{(g^{\prime},A^{\prime})}<\max\kappa$. Then we argue as in [16, Lemma 6.1]. ∎
5. Rigidity Theorems
In this section, we go in more detail into the case (2) of Theorem 4.1. Recall that case (2) of Theorem 1.3 involves a rather strong rigidity statement, as it implies that a metric on $M$ in class (2) with $R_{g}\geq 0$ is necessarily Ricci-flat. We will see that something similar happens in our situation as well.
The analogue of the Kazdan-Warner case (2) in our situation is the case of a simply connected non-spin spinc manifold $(M,L)$ with $\alpha^{c}(M,L)\neq 0$ (this of course means the Dirac operator $D_{L}$ has to have a nontrivial kernel) but with a pair $(g,A)$ for which $\mu_{1}(L_{(g,A)})=0$.
The simplest example of this situation is $M=\mathbb{C}\mathbb{P}^{2}$ with the Fubini-Study metric $g_{FS}$, $L=\mathcal{O}(3)$ (this is the smallest ample line bundle with odd Chern class and nonvanishing $\alpha^{c}(M,L)$, and is also the anti-canonical bundle), and the standard connection $A$ on $L$ with curvature $3\omega$, where $\omega$ is the Kähler form. Since the scalar curvature $R_{g_{FS}}=6$, then $R^{\text{{gen}}}_{(g_{FS},A)}=R_{g_{FS}}-2\cdot 3|\omega|_{op}\equiv 0$ (see also the calculation in the proof of [5, Corollary 5.2]). The next theorem shows that in some sense, this example is typical for this unusual situation.
Theorem 5.1.
Let $(M,L)$ be a simply connected spinc manifold which is non-spin, and assume that $\alpha^{c}(M,L)\neq 0$. Suppose there is a pair $(g,A)$ (a Riemannian metric $g$ on $M$ and a connection $A$ on $L$) for which $\mu_{1}(L_{(g,A)})=0$. Then $(M,g)$ is pointwise conformal to another metric $g^{\prime}$ for which $R^{\text{{gen}}}_{(g^{\prime},A)}$ vanishes identically. Moreover, then $(M,L)$ with metric $g^{\prime}$ admits a parallel spinor, and $(M,g^{\prime})$ is a Riemannian product of a Kähler manifold and a simply connected spin manifold with a parallel spinor. In particular, if $M$ does not split as a product, then $M$ is conformally Kähler, and $L$ is either the canonical or the anti-canonical line bundle on $M$.
Proof.
We apply the argument in the proof of Lemma 3.7. Let $\varphi$ be a strictly positive eigenfunction for the first eigenvalue $\mu_{1}(L_{(g,A)})=0$. Replace $g$ by the pointwise conformal metric $g^{\prime}=\varphi^{\frac{4}{n-2}}g$. The calculation in the proof of Lemma 3.7 gives that $R^{\text{{gen}}}_{(g^{\prime},A)}\equiv 0$, and the twisted conformal Laplacian for this new metric is just a multiple of the usual Laplacian on functions, so its lowest eigenvalue is of course still 00 (with eigenfunction $1$). Replacing $g$ by $g^{\prime}$, we now have $D_{L}^{2}=\nabla^{*}\nabla+R^{\text{{tw}}}_{(g^{\prime},A)}$, where $R^{\text{{tw}}}_{(g^{\prime},A)}\geq 0$ everywhere, with equality on the 00-eigenspace of $c(\Omega)$. Suppose $\psi$ is a harmonic spinor. (These exist since $\alpha^{c}(M,L)\neq 0$.) We have
$0=\langle D_{L}\psi,D_{L}\psi\rangle=\|\nabla\psi\|^{2}+\frac{1}{4}\langle R^{\text{{tw}}}_{g_{1},A}\psi,\psi\rangle,$
and since both terms on the right are nonnegative, they both have to vanish. Thus any harmonic spinor is necessarily parallel. The last statement follows from [21, Theorem 1.1]. ∎
Corollary 5.2.
Suppose $(M,L)$ is a simply connected spinc manifold which is not spin. Assume that $M$ does not split as a product, does not admit a Kähler metric, and satisfies $\alpha^{c}(M,L)\neq 0$. Then for any choice of $(g,A)$ on $(M,L)$, the lowest eigenvalue $\mu_{1}$ of the twisted conformal Laplacian must be negative.
Proof.
If $\mu_{1}(L_{(g,A)})$ were positive, then by Lemma 3.7, we could change the metric so that $D_{L}^{2}$ is strictly positive, which of course would imply that $\alpha^{c}(M,L)=0$, a contradiction. And if $\mu_{1}(L_{(g,A)})$ were 00, Theorem 5.1 would apply, which again would give a contradiction. Thus $\mu_{1}(L_{(g,A)})<0$. ∎
Example 5.3.
For example, suppose $M=\mathbb{C}\mathbb{P}^{2n}\#\mathbb{C}\mathbb{P}^{2n}$. This manifold does not split as a product and does not even admit an almost complex structure [10], so certainly it is not Kähler for any metric. The spinc line bundles $L$ on $M$ are parameterized by pairs $(k,\ell)$ of odd integers (representing the two components of $c_{1}(L)$), and
$\alpha^{c}(M,L)=\alpha^{c}(\mathbb{C}\mathbb{P}^{2n},\mathcal{O}(k))+\alpha^{c}(\mathbb{C}\mathbb{P}^{2n},\mathcal{O}(\ell))$
can only vanish if $|k|$ and $|\ell|$ are both less than $12n+$ 1 [12]. So if one or the other of $k$ and $\ell$ is at least $12n+$ 1 in absolute value, $\mu_{1}$ must be negative for any choice of $(g,A)$ on $(M,L)$.
Example 5.4.
Let $M$ be an Enriques surface, that is a complex surface double-covered by a K3 surface. Since a K3 surface has $\widehat{A}$-genus $2$, $M$ has $\widehat{A}$-genus $1$, which contradicts Rokhlin’s theorem, and thus $M$ cannot be spin. However, it is spinc, with a canonical spinc line bundle $L$ for which $c_{1}(L)$ is the unique non-zero class in $H^{2}(M,\mathbb{Z})_{\text{tors}}\cong\mathbb{Z}/2$. It is also well-known (by Yau’s theorem) that $M$ admits a Ricci-flat Kähler-Einstein metric $g$. Since $L$ admits a flat connection $A$ (since $L$ comes from the fundamental group of $M$), we have $R^{\text{{gen}}}_{(g,A)}\equiv 0$, and $M$ belongs to class (2) in Theorem 4.1. Of course, this $M$ is not simply connected.
We also observe that there is a partial converse to Theorem 5.1 that applies to Kähler-Einstein Fano varieties. Once again, this generalizes our previous example of $(\mathbb{C}\mathbb{P}^{2},\mathcal{O}(3))$. A very similar formulation appears in [11].
Theorem 5.5.
(cf. Goette-Semmelmann, [11, Theorem 1.9]) Let $(M,g)$ be a connected Kähler-Einstein manifold with positive Ricci curvature. Let $L$ be the anti-canonical line bundle on $M$. Then $(M,L)$ does not admit $(g,A)$ with $R^{\text{{gen}}}_{(g,A)}>0$, but with the metric $g$ and connection $A$ on $L$ with harmonic curvature (necessarily a multiple of $g$) $R^{\text{{gen}}}_{(g,A)}=0$.
Proof.
Note that $\alpha^{c}(M,L)$ for this choice of $L$ coincides with the Todd genus, which is the alternating sum of the Hodge numbers $h^{0,q}(M)$. Since the Ricci curvature is positive, the $h^{0,q}$ vanish for $q>0$ by the Kodaira vanishing theorem, the Bochner formula, and Serre duality, so the Todd genus is $1$ [19, Corollary IV.11.12]. Thus $M$ cannot admit positive generalized scalar curvature. On the other hand, since $M$ is Kähler-Einstein, the curvature $2$-form of $L$ (with its standard connection) coincides with the Ricci form, and the Ricci tensor is $\frac{R_{g}}{2}g$. That implies that $2|\Omega|_{op,g}=R_{g}$, so that $R^{\text{{gen}}}_{g}=0$. (See [19, Lemma IV.11.10] for a similar calculation.) ∎
Remark 5.6.
While Theorem 5.5 and [11, Theorem 1.9] both address the nonexistence of metrics with scalar curvature bounded below by curvature form norms, it is important to note that the specific norms used in each theorem are distinct.
6. The Index Difference Map
We start with some basic definitions.
Definition 6.1.
Let $M$ be a closed spin manifold, $\dim M=n$, with a fixed psc metric $g_{0}$. The index difference homomorphism
$\operatorname{inddiff}(M)\colon\,\pi_{q}(\mathcal{R}^{+}(M))\to KO_{q+n+1}$
is defined by sending the homotopy class of $f\colon\,(S^{q},*)\to(\mathcal{R}^{+}(M),g_{0})$ to the index obstruction to filling in $f$ to a map $(D^{q+1},*)\to(\mathcal{R}^{+}(M),g_{0})$, or in other words the $KO_{*}$-valued index of the Dirac operator on the spin manifold $N^{n+q+1}=M^{n}\times D^{q+1}\cup_{M^{n}\times S^{q}\times\{0\}}M^{n}\times S^{q}\times[0,\infty)$ with Riemannian metric restricting to $f(x)+dt^{2}$ on $M\times\{x\}\times[0,\infty)$, for $x\in S^{q}$.
The index above is well-defined because the scalar curvature of $N$ is uniformly positive outside of a compact set, and thus the Dirac operator on $N$ is Fredholm; see [7, 8] for detailed descriptions of the index difference.
We can do something analogous in the spinc case.
Definition 6.2.
Let $(M,L)$ be a spinc manifold with a fixed choice of basepoint $(g_{0},A_{0})\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$. The spinc index difference map
$\operatorname{inddiff}^{c}(M,L)\colon\,\pi_{q}(\mathcal{R}^{{\mathrm{gen}+}}(M,L))\to KU_{q+n+1}$
is defined just as in the spin case, by sending the homotopy class of
$f\colon\,(S^{q},*)\to(\mathcal{R}^{{\mathrm{gen}+}}(M,L),(g_{0},A_{0}))$
to the index obstruction to filling in $f$ to a map $(D^{q+1},*)\to(\mathcal{R}^{{\mathrm{gen}+}}(M,L),(g_{0},A_{0}))$, by taking the usual $KU_{*}$-valued index of the complex Dirac operator $D_{\tilde{L}}$ on the spinc manifold $(N^{n+q+1},\tilde{L}):=(M^{n}\times D^{q+1}\cup_{M^{n}\times S^{q}\times\{0\}}M^{n}\times S^{q}\times[0,\infty),\tilde{L})$, with the same condition on a Riemannian metric as in the spin case, and with $\tilde{L}=p^{*}L$, where $p$ is the natural projection $p:N^{n+q+1}\to M$.
The following propositions show that the index difference map is nontrivial (Theorem 1.9, (i)).
Proposition 6.3.
Let $(M,L)$ be a closed connected $n$-dimensional spinc manifold, with $n=4k-1$, $k\geq 2$ and admitting generalized positive scalar curvature. Then the map
$\operatorname{inddiff}^{c}(M,L)\colon\,\pi_{0}(\mathcal{R}^{{\mathrm{gen}+}}(M,L))\to\mathbb{Z}$
takes infinitely many nonzero values, and in particular, $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ has infinitely many connected components.
Proof.
We choose a $2$-connected spin manifold $N^{4k}$ with $\widehat{A}(N)\neq 1$. If $k$ is even, this can be a product of copies of the “Bott manifold,” for which a construction is given in [6]. If $k\geq 2$ is odd, $\widehat{A}(N)$ is necessarily even and one can construct $N$ as follows. Use an $E_{8}$ plumbing [2] to get a parallelizable $4k$-manifold $Z^{4k}$ with boundary a homotopy sphere $\Sigma^{4k-1}$ (a generator of $bP_{4k}$). This homotopy sphere has finite order in the group $\Theta_{4k-1}$ of homotopy spheres, so we can take a boundary connected sum of finitely many copies of $Z^{4k}$ to get a parallelizable manifold $V^{4k}=Z^{4k}\#_{\partial}\cdots\#_{\partial}Z^{4k}$ with boundary $S^{4k-1}$. Glue in a disk:
$N^{4k}:=V^{4k}\cup_{S^{4k-1}}D^{4k}.$
Now $N^{4k}$ is our desired manifold: it is closed $4k$-manifold which is highly connected, hence spin, and has only one nonzero Pontryagin class, the top Pontryagin class $p_{k}$. Since the signature of $Z^{4k}$ is $8$, $N$ has nonzero signature and nonzero $\widehat{A}$-genus.
Now we turn to our spinc manifold $(M,L)$. We choose $(g,h,A)\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ and consider the spinc bordism $W=(M\times[0,1])\#N$ from $M$ to itself, where $N$ is attached to the interior of the cylinder $M\times[0,1]$. The bordism $W$ has natural spinc structure coming from $M$, since the spinc structure on $M$ defines one on the cylinder $M\times[0,1]$ and then this spinc structure extends uniquely to $W$ since $N$ is $2$-connected. We use a product metric $g+dt^{2}$ to extend the metric from $M\times\{0\}$ to $M\times[0,\varepsilon]$ for some $\varepsilon>0$. By the proof of the bordism theorem [5, Theorem 4.3] and the fact that the inclusion of either boundary copy of $M$ into $W$ is $2$-connected, we can push the triple $(g,h,A)$ across the bordism $W$ to obtain $(g^{\prime},h^{\prime},A^{\prime})\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ at $M\times\{1\}$. The value of the index difference is then obtained by computing the index of the Dirac operator on the manifold
${W^{\prime}:=}M\times(-\infty,0]\cup_{M\times\{0\}}W\cup_{M\times\{1\}}M\times[1,\infty)$
using $(g+dt^{2},h,A)$ on $M\times(-\infty,0]$ and $(g^{\prime}+dt^{2},h^{\prime},A^{\prime})$ on $M\times[1,\infty)$. This is just the sum of the index of Dirac on $M\times\mathbb{R}$, which vanishes, with $\widehat{A}(N)$,555 This is because there is bordism, just a product with $[0,1]$ off a compact set, between our manifold $W^{\prime}$ and $(M\times\mathbb{R},g+dt^{2})\sqcup N$, for which the index is obviously the sum of indices of the disjoint manifolds, and because of the bordism invariance of the index of the Dirac operator. and so the index difference is nontrivial. We can now get infinitely many different values for the index difference by replacing $N$ by $\overbrace{N\#N\#\cdots\#N}^{k}$ in this construction to get an index difference of $k\widehat{A}(N)$ (for any positive $k$). ∎
Before stating the next result, we need some notation. Let $c\colon\,KO_{n}\to KU_{n}$ denote the “complexification map” (which is zero if $n$ is not divisible by $4$, an isomorphism if $n\equiv 0\pmod{8}$, and a monomorphism with cokernel $\mathbb{Z}/2$ if $n\equiv 4\pmod{8}$). We also need an observation: if $n$ is divisible by 4, there is always a closed $n$-dimensional spinc manifold $(M,L)$ which is not spin with $\alpha^{c}(M,L)=1$. For example, we can take $M=\mathbb{C}\mathbb{P}^{n/2}$ and $L=\mathcal{O}\bigl{(}1+\frac{n}{2}\bigr{)}$. Take such an $(M,L)$ and fix a pair $(g_{M},A)$, which cannot have generalized scalar curvature $R^{\text{{gen}}}_{(g,A)}$ positive since the index of $D_{L}$ is nonzero.
Consider the spinc manifold $(S^{k}\times M,p^{*}L)$, where $k\geq 2$, and let $p\colon\,S^{k}\times M\to M$ be projection on the second factor. The connection $A$ on $L$ gives a connection $p^{*}A$ on $p^{*}L$. Let $\mathcal{R}^{+}(S^{k})$ be the space of psc metrics on $S^{k}$, with basepoint the standard round metric. Now we define a map
(6.1) $\sigma\colon\,\pi_{q}(\mathcal{R}^{+}(S^{k}))\to\pi_{q}(\mathcal{R}^{{\mathrm{gen}+}}(S^{k}\times M))$
as follows. Fix a map $f\colon\,S^{q}\to\mathcal{R}^{+}(S^{k})$ representing a homotopy class, and note that we can find a scale $t>0$ small enough so that the generalized scalar curvature
$R^{\text{{gen}}}_{(t^{2}g+g_{M},p^{*}A)}=(t^{-2}R_{g}+R_{g_{M}})-2|\Omega|_{op}$
is positive for all $g\in f(S^{q})\subset\mathcal{R}^{+}(S^{k})$, the compact image of our map $S^{q}\to\mathcal{R}^{+}(S^{k})$. Then the formula $x\mapsto(t^{2}f(x)+g_{M},p^{*}A)$, $x\in S^{q}$, defines $\sigma([f])$, since it’s clear that up the homotopy class of this family of metrics only depends on the homotopy class of $f$ and not on the parameter $t$. Once $t$ is fixed, $\sigma$ is obtained by taking the product of the family of metrics $t^{2}f$ (which represents the same homotopy class as $f$) with the metric $g_{M}$ on $M$, and this operation of direct product is compatible with addition of homotopy classes, so in this way we get a homomorphism $\pi_{q}(\mathcal{R}^{+}(S^{k}))\to\pi_{q}(\mathcal{R}^{{\mathrm{gen}+}}(S^{k}\times M))$. We obtain the following diagram:
[…]
where $\tilde{c}=\mathbb{B}\circ c$, and $\mathbb{B}$ is the Bott isomorphism.
Proposition 6.4.
Let $(M,L)$ be as above and $q+k+1$ be a multiple of 4. Then the diagram (6.2) commutes.
Proof.
Suppose we are given a map $g\colon\,(S^{q},*)\to(\mathcal{R}^{+}(S^{k}),g_{0})$, where $g_{0}$ is the standard metric. By scaling $g$ we obtain a map $\sigma(g)\colon\,(S^{q},*)\to(\mathcal{R}^{{\mathrm{gen}+}}(S^{k}\times M),g_{0}+g_{M})$. Then $\operatorname{inddiff}(g)$ is defined by extending $g$ from a family of metrics on $S^{q}\times S^{k}$ to a family of metrics on $D^{q+1}\times S^{k}$ which are product metrics on a collar neighborhood of the boundary, and for each $x\in S^{q}$, extending via a cylinder metric $g(x)+dt^{2}$ on $S^{k}\times[0,\infty)$ to get a complete metric on $\mathbb{R}^{q+1}\times S^{k}$ with positive scalar curvature off a compact set. The complex Dirac operator for this metric is Fredholm, and has index given by $c(\operatorname{inddiff}(g))$. Now take the Riemannian product with $(M,g_{M})$ and equip this with the line bundle $p^{*}L$. We can repeat the same construction as before with the spinc Dirac operator, and since this is the external product of the spinc Dirac on $M$ with the Dirac on $D^{q+1}\times S^{k}$, the indices multiply via the multiplicativity of the index as explained in [1, §5 and §9]; here we use that $\alpha^{c}(M,L)=1$. So the conclusion follows. ∎
Now we can prove (ii) from Theorem 1.9.
Corollary 6.5.
There exists a simply connected spinc manifold $(M,L)$, not spin, for which infinitely many homotopy groups of $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ are non-zero.
Proof of Corollary.
Simply take $M$ to be $\mathbb{C}\mathbb{P}^{2}\times S^{7}$, $L=p^{*}\mathcal{O}(3)$, $p$ the projection of $\mathbb{C}\mathbb{P}^{2}\times S^{7}$ onto $\mathbb{C}\mathbb{P}^{2}$. If $x$ is the standard generator of $H^{2}(\mathbb{C}\mathbb{P}^{2};\mathbb{Z})$, the $\widehat{A}$ class is $\widehat{\mathbf{A}}(\mathbb{C}\mathbb{P}^{2})=1-\frac{x^{2}}{8}$ and so $\operatorname{ind}D_{\mathcal{O}(3)}$ is the coefficient of $x^{2}$ in $\left(1-\frac{x^{2}}{8}\right)(1+\frac{3x}{2}+\frac{9x^{2}}{8})$, which is $1$, and use the known nontriviality of the index difference map for $S^{7}$ (see for example [4, Theorem A]). ∎
Remark 6.6.
We stated in Conjecture 1.7 that nontriviality of the topology of $\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ is in fact quite a general phenomenon. So one shouldn’t attach much significance to the special nature of this example.